# Intra-Trajectory Consistency for Reward Modeling

### v2

@@ 4.1 Experimental Setup
Training Details. (1) We train the proposed reward modeling method based on GRM (Yang et al., 2024a), which achieves competitive results in the RewardBench benchmark. The resulting reward model is termed ICRM. We validate the proposed method on Gemma-2B-it (Mesnard et al., 2024), Llama3-8B-instruct (Dubey et al., 2024), and Qwen-1.5B-Instruct. The hyperparameter $\alpha$ is set to 0.1 in all of our experiments. Additional training configurations, including the procedure for training Qwen-1.5B-Instruct with unpaired data, are detailed in Appendix B. (2) For RLHF, we adopt DPO strategy (Rafailov et al., 2023) to train the policy model with Gemma-2B-it as the initial model. Following (Dong et al., 2024; Liu et al., 2024), we generate 8 responses for each prompt in the training set of RLHF. Then these responses are scored by the 2B reward model trained with Unified-Feedback 400K. The best-worst response pairs for each prompt are used to train the DPO policy. All experiments are implemented on at most two A800 GPUs.
@@ Appendix A Related Works
Generalization of Reward Models. The generalization of reward models to unseen responses is essential for improving their robustness in RLHF and inference-time verification (Gao et al., 2023; Yang et al., 2024a). To improve it, multiple approaches have been developed, such as ensemble techniques (Eisenstein et al., 2023; Rame et al., 2024) , data augmentation (Shen et al., 2024; Liu et al., 2024), direct correction of measurable bias (Dubois et al., 2024; Chen et al., 2024a), and hidden-state regularization (Yang et al., 2024a). For example, (Coste et al., 2024) proposes to learn multiple estimators and combine them to improve the robustness of the rewards. (Liu et al., 2024) introduces a data augmentation approach derived from the causal framework to differentiate between contextual signals and context-free artifacts. (Chen et al., 2024a) proposes a framework trained to predict both rewards and lengths so that it can disentangle the representation of the content quality from the lengths of responses. (Yang et al., 2024a) introduces SFT and DPO losses to regularize the hidden states of reward models. While these methods strengthen RMs for AI alignment, their effectiveness remains constrained by sparse response-level supervision, limiting further generalization.
Outcome Rewards and Process Rewards. Outcome rewards and process rewards represent two fundamental reward paradigms in reinforcement learning and AI alignment (Schulman et al., 2017; Im and Li, 2024; Uesato et al., 2022). Outcome rewards, which evaluate final results (e.g., task completion), are simple to define and require minimal annotation effort. However, they often fail to guide desirable behaviors in complex, long-horizon tasks (Lightman et al., 2023). In contrast, process rewards, particularly critical in autoregressive generation, capture causal dependencies between tokens by evaluating intermediate steps (e.g., reasoning coherence), as demonstrated by (Li and Li, 2025). While effective, process reward learning typically demands costly step-level annotations (Wang et al., 2024a; Luo et al., 2024). Recent work mitigates this through automated methods, such as (Wang et al., 2024a)’s stepwise scoring based on downstream answer correctness. Yet, such approaches remain suboptimal without a free annotator, leaving response-level annotation the prevailing standard for reward modeling. Therefore, it is meaningful to explore methods for incorporating more process-related signals based solely on response-level labeling.
@@ Appendix B Implementation Details
Overall Information. Our code is based on LlamaFactory (Zheng et al., 2024), a powerful code framework that supports multiple LLMs training strategies. The version number we adopted is 0.9.1.dev0. The LLMs involved in the experiments are all from HuggingFace. In the experiments, the models are trained based on LoRA (Hu et al., 2022). We employ the DeepSpeed framework [^4] https://github.com/deepspeedai/DeepSpeed?tab=readme-ov-file to enhance training efficiency and reduce GPU memory requirements. For models less than 2B parameters, we utilize the ZeRO-2 offload configuration. When training includes 8B models, we adopt the ZeRO-3 offload setting.
More Details for Training of Reward Models. Our reward model training procedure primarily follows GRM (Yang et al., 2024a), specifically adopting their GRM-sft configuration. GRM-sft incorporates an additional SFT loss alongside the standard reward loss to regularize hidden-state representations, with a default weighting factor of 0.001. The reward model is initialized using the backbone of a pre-trained LLM with an additional randomly initialized linear logistic regression layer (dropout=0.1). Common hyper-parameters are detailed in Table 6. Key variations include batch size and training epochs: for datasets less than 40K, we use batch size 12 with 2 epochs; other experiments employ batch size 24 with 1 epoch. Notably, our proposed method occasionally requires a generator during reward model training. The generator for the 2B model is trained with 40K samples from the Unified-Feedback dataset. The generator for the 8B model is trained with 40K samples from the Unified-Feedback dataset and the Skywork dataset. When training, it shares the same configuration as the reward model (Table 6), except for a reduced learning rate (1e-7) and a single epoch to prevent overfitting. For the reward model that is only trained with the Unified-Feedback dataset, we set the sequence cutoff length to 1,024 tokens. All other reward models and generators use a 2,048-token cutoff to better handle long-form data.
@@ C.1 Results on Different Sizes of Training Samples
We further investigate the performance of the proposed method on different training data scales, with results presented in Table 7. The experimental results demonstrate that the reward model trained with the proposed method consistently outperforms the baseline GRM method, validating its stability across varying amounts of training data.

### v3

@@ 4.1 Experimental Setup
Training Details. (1) We train the proposed reward modeling method based on GRM (Yang et al., 2024a), which achieves competitive results in the RewardBench benchmark. The resulting reward model is termed ICRM. We validate the proposed method on Gemma-2B-it (Mesnard et al., 2024), Llama3-8B-instruct (Dubey et al., 2024), and Qwen-1.5B-Instruct. The hyperparameter $\alpha$ is set to 0.1 in all of our experiments. Additional training configurations, including the procedure for training Qwen-1.5B-Instruct with unpaired data, are detailed in Appendix B. (2) For RLHF, we adopt DPO strategy (Rafailov et al., 2023) to train the policy model with Gemma-2B-it as the initial model. Following (Dong et al., 2024; Liu et al., 2024), we generate 8 responses for each prompt in the training set of RLHF. Then these responses are scored by the 2B reward model trained with Unified-Feedback 400K. The best-worst response pairs for each prompt are used to train the DPO policy. All experiments are implemented on at most two A800 GPUs, each with 80GB of memory.
@@ 4.3 Evaluation on RLHF
4.4 Evaluation on Inference-Time Verification
4.5 Empirical Analysis
@@ Appendix A Related Works
Generalization of Reward Models. The generalization of reward models to unseen responses is essential for improving their robustness in RLHF and inference-time verification [Gao et al., 2023, Yang et al., 2024a]. To improve it, multiple approaches have been developed, such as ensemble techniques [Eisenstein et al., 2023, Rame et al., 2024] , data augmentation [Shen et al., 2024, Liu et al., 2024], direct correction of measurable bias [Dubois et al., 2024, Chen et al., 2024a], and hidden-state regularization [Yang et al., 2024a]. For example, [Coste et al., 2024] proposes to learn multiple estimators and combine them to improve the robustness of the rewards. [Liu et al., 2024] introduces a data augmentation approach derived from the causal framework to differentiate between contextual signals and context-free artifacts. [Chen et al., 2024a] proposes a framework trained to predict both rewards and lengths so that it can disentangle the representation of the content quality from the lengths of responses. [Yang et al., 2024a] introduces SFT and DPO losses to regularize the hidden states of reward models. While these methods strengthen RMs for AI alignment, their effectiveness remains constrained by sparse response-level supervision, limiting further generalization.
Outcome Rewards and Process Rewards. Outcome rewards and process rewards represent two fundamental reward paradigms in reinforcement learning and AI alignment [Schulman et al., 2017, Im and Li, 2024, Uesato et al., 2022]. Outcome rewards, which evaluate final results (e.g., task completion), are simple to define and require minimal annotation effort. However, they often fail to guide desirable behaviors in complex, long-horizon tasks [Lightman et al., 2023]. In contrast, process rewards, particularly critical in autoregressive generation, capture causal dependencies between tokens by evaluating intermediate steps (e.g., reasoning coherence), as demonstrated by [Li and Li, 2025]. While effective, process reward learning typically demands costly step-level annotations [Wang et al., 2024a, Luo et al., 2024]. Recent work mitigates this through automated methods, such as [Wang et al., 2024a]’s stepwise scoring based on downstream answer correctness. Yet, such approaches remain suboptimal without a free annotator, leaving response-level annotation the prevailing standard for reward modeling. Therefore, it is meaningful to explore methods for incorporating more process-related signals based solely on response-level labeling.
@@ Appendix B Implementation Details
Overall Information. Our code is based on LlamaFactory [Zheng et al., 2024], a powerful code framework that supports multiple LLMs training strategies. The version number we adopted is 0.9.1.dev0. The LLMs involved in the experiments are all from HuggingFace. In the experiments, the models are trained based on LoRA [Hu et al., 2022]. We employ the DeepSpeed framework [^4] https://github.com/deepspeedai/DeepSpeed?tab=readme-ov-file to enhance training efficiency and reduce GPU memory requirements. For models less than 2B parameters, we utilize the ZeRO-2 offload configuration. When training includes 8B models, we adopt the ZeRO-3 offload setting.
More Details for Training of Reward Models. Our reward model training procedure primarily follows GRM [Yang et al., 2024a], specifically adopting their GRM-sft configuration. GRM-sft incorporates an additional SFT loss alongside the standard reward loss to regularize hidden-state representations, with a default weighting factor of 0.001. The reward model is initialized using the backbone of a pre-trained LLM with an additional randomly initialized linear logistic regression layer (dropout=0.1). Common hyper-parameters are detailed in Table 6. Key variations include batch size and training epochs: for datasets less than 40K, we use batch size 12 with 2 epochs; other experiments employ batch size 24 with 1 epoch. Notably, our proposed method occasionally requires a generator during reward model training. The generator for the 2B model is trained with 40K samples from the Unified-Feedback dataset. The generator for the 8B model is trained with 40K samples from the Unified-Feedback dataset and the Skywork dataset. When training, it shares the same configuration as the reward model (Table 6), except for a reduced learning rate (1e-7) and a single epoch to prevent overfitting. For the reward model that is only trained with the Unified-Feedback dataset, we set the sequence cutoff length to 1,024 tokens. All other reward models and generators use a 2,048-token cutoff to better handle long-form data.
@@ C.1 Results on Different Sizes of Training Samples
We further investigate the performance of the proposed method on different training data scales. We show the accuracy results on RewardBench with different sizes of training samples from the Unified-Feedback dataset in Table 7. The experimental results demonstrate that the reward model trained with the proposed method consistently outperforms the baseline GRM method, validating its stability across varying amounts of training data.